QUANTITATIVE ABSORPTION MEASUREMENTS USING CAVITY RING DOWN SPECTROSCOPY

\underline{R.D. van Zee}, J.P. Looney, J.T. Hodges, Process Measurements Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

This paper investigates the effects of cavity design and laser bandwidth on quantitative cavity ring-down spectroscopy using the ^{r}R transitions of the X(v = 0) to b(v = 0) band of molecular oxygen. It is found that failure to account for laser bandwidth leads to systematic errors in the number densities determined from measured ring-down signals. By fitting the frequency integrated expression for the ring-down signal and using the measured laser lineshapes, excellent agreement between measured and and predicted number densities is found. The quantitative measurement of gas densities over a wide dynamic range will be presented.